OpenStudy (anonymous):

find an equation of the line passing through the point (2,3) and parallel to the line 3x+4y-8=0

OpenStudy (anonymous):

given point (2,3) and the line 3x+4y-8=0 first we want to simplify 3x+4y-8=0 to 4y= -3x + 8 y = -(3/4)x + 2 since this is parallel to the line we wanted its slope m is the same whats the slope m= ?

OpenStudy (anonymous):

the slope is -3/4

OpenStudy (anonymous):

if thats the slope we can use it to form equation of the line using the formula and the point (2,3) y-y1=m(x - x1) try forming it using this formula

OpenStudy (anonymous):

y-3=-3/4(x-2)

OpenStudy (anonymous):

yesss... now you can rearrange your answer to slope intercept form

OpenStudy (anonymous):

that is where I get lost

OpenStudy (anonymous):

ah ok, try to rearrange it to slope intercept form

OpenStudy (anonymous):

so everything is alright now? :D have fun solving now... :D